OpenStudy (1018):

a conical container has a height of 20cm and a base diameter of 12cm. It holds water a depth of 10cm above its vertex. Find the volume of its content.

OpenStudy (dayakar):

height (h)=20cm redius(r) = diameter/2 =12/2 =6cm depth(h) = 10cm

OpenStudy (dayakar):

\[volume(v) =\frac{ 1 }{ 3 }*\pi* r ^{2}*h\]

OpenStudy (dayakar):

substitute above values u get the answer

OpenStudy (1018):

hey sorry for the late reply. I tried to substitute both 'h' in the equation but the answers weren't in the choices.